# Dissipation Pathways in a Photosynthetic Complex

### v1

@@ Conclusions
The ability to precisely map dissipation pathways in complex chemical systems, as demonstrated in this study of the FMO complex, opens new avenues for investigating the role of the environment in energy transfer dynamics. We envision future applications of this methodology extending beyond the FMO complex to other light-harvesting systems, including, but not limited to, light-harvesting complex II (LH2)Mirkovic et al. (2017), phycobiliprotein PC645Blau et al. (2018), cyanobacterial light-harvesting proteinsWomick and Moran (2011) and other molecular systemsLee et al. (2017); Novoderezhkin et al. (2004); Gustin et al. (2023); Rätsep and Freiberg (2007); Wiethorn et al. (2023). Furthermore, our strategy can aid in guiding emerging efforts to develop more efficient artificial light-harvesting systemsHäse et al. (2020) by identifying the key vibrational modes and pigment molecules involved in energy transfer in natural light-harvesting systems and, more broadly, to advance efforts to control quantum systems via environment engineering.
Materials and Methods
@@
Here, we outline the main ideas of our quantum theory of dissipation pathways based on Markovian quantum master equations describing the open quantum dynamics. We focus on the case when the environment consists of a local set of effective harmonic modes, as is the case with photosynthetic complexes. However, the approach can be applied to any harmonic or anharmonic environment, provided that it can be modeled with independent environment degrees of freedom.Kim and Franco (2024a, 2021)
We consider an array of $N$ chromophores that are coupled to one another and to local thermal environments. The state |$\ket{A}$ describes a physical situation in which the chromophore |$\ket{A}$ is in the excited electronic state while the other chromophores in the complex are in the electronic ground state. The population dynamics of a chromophore |$\ket{A}$ is given by
@@
where $K_{BA}$ is the rate constant of excitation transfer from |$\ket{A}$ to |$\ket{B}$ given by
@@
×$\times\exp\left[-g_{A}(t^{\prime})-g_{B}(t^{\prime})\right]dt^{\prime}.$
Here, $E_{A}$ and $E_{B}$ represent the electronic energy gaps from the ground to the excited state at chromophores |$\ket{A}$ and |$\ket{B}$, respectively, and $V_{AB}$ the electronic coupling between these chromophores. The total electron-vibrational coupling at a given chromophore |$\ket{A}$ is measured by the reorganization energy $\Lambda_{A}$. The reorganization energy is related to the environment spectral density by
@@
Within this framework, the dissipation rate $\dot{E}_{Aj}(t)$ in the $j$-th vibrational mode belonging to the chromophore |$\ket{A}$ can be evaluated as
@@
Here, $\lambda_{A}^{j}$ is the discrete reorganization energy of the $j$-th vibration that belongs to |$\ket{A}$. If we sum over all the vibrations at chromophore |$\ket{A}$ we obtain the total reorganization energy $\Lambda_{A}=\sum_{j}\lambda_{A}^{j}$. In turn, $\mathcal{I}_{BA}(\omega)$ is the “dissipative potential” defined as
@@
×$\times\exp\left[-g_{A}(t^{\prime})-g_{B}(t^{\prime})\right]$
×$\times\left[\cos(\omega t^{\prime})-i\coth\left(\frac{\beta\hbar\omega}{2}\right)\sin(\omega t^{\prime})\right]dt^{\prime}$
which quantifies the ability of the $j$-th vibration to induce dissipation per unit of reorganization energy at chromophore |$\ket{A}$. To obtain an expression for the rate of dissipation at chromophore |$\ket{A}$ as a function of frequency at a certain time, $\mathcal{D}_{A}(\omega,t)$, we replace $\mathcal{K}_{BA}^{Aj}$ and $\mathcal{K}_{AB}^{Aj}$ by the “dissipative spectral densities” $\mathcal{J}_{BA}^{A}(\omega)$ and $\mathcal{J}_{AB}^{A}(\omega)$
@@
This equation enables us to compute the dissipation of chromophore |$\ket{A}$ as a function of time and environment frequency. Importantly, all the terms in Eq. (9) can be obtained without explicitly accessing the dynamics of the vibrational environment, keeping the computational cost tractable.
@@
[…]
@@
Non-Markovian Effects. Equation (1) is a Markovian quantum master equation which assumes that the dynamics of the environment is fast compared to the system’s dynamics. In our simulations, we account for non-Markovian effects that go beyond this treatment by using a time scale separation methodBerkelbach et al. (2012); Montoya-Castillo et al. (2015) to split the spectral density, $J(\omega)$, into “slow” and “fast” components, denoted as $J_{\text{slow}}(\omega)$ and $J_{\text{fast}}(\omega)$, respectively. To introduce the desired non-Markovianity, we prevent the environment modes in $J_{\text{slow}}(\omega)$ from directly influencing the system dynamics. The spectral density separation is formally achieved by defining
@@
Next, we treat $J_{\text{slow}}(\omega)$ as a source of static disorder that modulates the Bchl’s electronic energies through a Gaussian random noise process with a standard deviation
[…]
@@
To account for the static disorder, we performed an ensemble average of over 10,000 independent noise realizations. Importantly, while each realization featured distinct site energies ($E$) drawn from a distribution representing static disorder, Eq. (14) , the underlying set of fast spectral density components ($J_{\text{fast}}(\omega)$) remained the same across all realizations. This significantly reduces the computational burden, as several crucial quantities, including the line-broadening functions ($g(t)$), total reorganization energies ($\Lambda$), and the discretized reorganization energies ($\lambda^{j}$ for each $j$-th vibrational mode), needed to be computed only once. Furthermore, to enhance computational efficiency, the calculation of the dissipative potential Eq. 7 was vectorized with respect to both the frequency ($\omega$) and time ($t$) grids using the NumPy library in Python. This vectorization allows for the simultaneous evaluation of these quantities for all frequency modes and time points, thereby significantly reducing the overall computation time.
[…]
Acknowledgements.

### v2

@@ Abstract
The ability to precisely map dissipation pathways in complex chemical systems, as demonstrated in this study of the FMO complex, opens new avenues for investigating the role of the environment in energy transfer dynamics. We envision future applications of this methodology extending beyond the FMO complex to other light-harvesting systems, including, but not limited to, light-harvesting complex II (LH2)mirkovic2017light, phycobiliprotein PC645blau2018local, cyanobacterial light-harvesting proteinswomick2011vibronic and other molecular systemslee2017first, novoderezhkin2004energy, gustin2023mapping, ratsep2007electron, wiethorn2023beyond. Furthermore, our strategy can aid in guiding emerging efforts to develop more efficient artificial light-harvesting systemshase2020designing by identifying the key vibrational modes and pigment molecules involved in energy transfer in natural light-harvesting systems and, more broadly, to advance efforts to control quantum systems via environment engineering.
1 Theoretical Methods
In Ref. kim2024general1 we developed MQME-D, a general framework for constructing practical and accurate schemes to isolate dissipation pathways in open quantum systems targeting Markovian Quantum Master Equations (MQME). The formulation incorporates a specific bath degree of freedom into the subsystem component and calculates the change in its energy with Nakajima-Zwanzig projection operator technique. By perturbatively expanding the Liouville-von Neumann equation, the MQME-D elucidates the rate of dissipation expressed by using traces of operator products, which can be applied for general types of bath and subsystem-bath interaction. The MQME-D also rigorously satisfies thermodynamic principles such as energy conservation and detailed balance.
@@ 1 Theoretical Methods
Here, we outline the main ideas of our quantum theory of dissipation pathways based on Fermi’s Golden Rule Markovian quantum master equations describing the open quantum dynamics. We focus on the case when the environment consists of a local set of effective harmonic modes, as is the case with photosynthetic complexes. However, the approach can be applied to any harmonic or anharmonic environment, provided that it can be modeled with independent environment degrees of freedom.kim2024general1, kim2021theory
[…]
@@
[…]
@@
$\times\exp\left[-g_{A}(t^{\prime})-g_{B}(t^{\prime})\right]dt^{\prime}.$
@@
$\Lambda_{A}=\int_{0}^{\infty}\frac{J_{A}(\omega)}{\omega}\>d\omega.$ (3)
@@
[…]
@@
[…]
@@
$\times\exp\left[-g_{A}(t^{\prime})-g_{B}(t^{\prime})\right]$
$\times\left[\cos(\omega t^{\prime})-i\coth\left(\frac{\beta\hbar\omega}{2}\right)\sin(\omega t^{\prime})\right]dt^{\prime}$
[…]
@@
where we have used the property $\sum_{j}\lambda_{A}^{j}=\int_{0}^{\infty}d\omega\>J_{A}(\omega)/\omega$. This replacement yields the dissipation function
@@
$\mathcal{E}_{A}(\omega,t)=\int_{0}^{t}\mathcal{D}_{A}(\omega,t^{\prime})dt^{\prime}.$ (10)
@@
Non-Markovian Effects. Equation (1) is a Markovian quantum master equation which assumes that the dynamics of the environment is fast compared to the system’s dynamics. In our simulations, we account for non-Markovian effects that go beyond this treatment by using a time scale separation methodberkelbach2012reduced, montoya2015extending to split the spectral density, $J(\omega)$, into “slow” and “fast” components, denoted as $J_{\text{slow}}(\omega)$ and $J_{\text{fast}}(\omega)$, respectively. To introduce the desired non-Markovianity, we prevent the environment modes in $J_{\text{slow}}(\omega)$ from directly influencing the system dynamics. The spectral density separation is formally achieved by defining
@@
Next, we treat $J_{\text{slow}}(\omega)$ as a source of quasi-static noise that modulates the Bchl’s electronic energies through a Gaussian random noise process with a standard deviation
$\sigma_{\text{slow}}=\frac{1}{\hbar}\int_{0}^{\infty}J_{\text{slow}}(\omega)\coth\left(\frac{\beta\hbar\omega}{2}\right)d\omega$ (14)
@@
To account for the non-Markovian effects, we performed an ensemble average of over 10,000 independent noise realizations. Importantly, while each realization featured distinct site energies ($E$) drawn from a distribution representing slow-bath component, Eq. (14) , the underlying set of fast spectral density components ($J_{\text{fast}}(\omega)$) remained the same across all realizations. This significantly reduces the computational burden, as several crucial quantities, including the line-broadening functions ($g(t)$), total reorganization energies ($\Lambda$), and the discretized reorganization energies ($\lambda^{j}$ for each $j$-th vibrational mode), needed to be computed only once. Furthermore, to enhance computational efficiency, the calculation of the dissipative potential Eq. 7 was vectorized with respect to both the frequency ($\omega$) and time ($t$) grids using the NumPy library in Python. This vectorization allows for the simultaneous evaluation of these quantities for all frequency modes and time points, thereby significantly reducing the overall computation time.
The fast spectral density components, $J_{\text{fast}}(\omega)$, were discretized using 4000 effective harmonic oscillator modes equally spaced in frequency. The time integrals required for calculating the population transfer rate constants (Eq. 2) and the dissipation rates (Eq. 8) were evaluated numerically using the trapezoidal rule. This method was applied over a finely spaced time grid, with a step size of $\Delta t=0.5\,\text{fs}$ and extending to a maximum integration time of $T_{\text{max}}=30\,\text{ps}$. The trapezoidal rule provides a good balance between accuracy and computational cost for these types of integrals. The propagation of the rate equations governing the electronic populations (Eq. 1) and the dissipation dynamics (Eq. 6) was performed using the fourth-order Runge-Kutta (RK4) method. A consistent time step of $0.5\,\text{fs}$ was employed for the RK4 integration, ensuring numerical stability and accuracy throughout the simulation. The accuracy of this approach in generating semi-quantitative results for the FMO complex is assessed against the numerically exact HEOM using illustrative models (see Secs. SII-SIII of the Supplementary Information). The codes used to obtain the results of this paper are available on GitHub.GustinGitHub
@@
The Supporting Information provides details regarding the applicability of the dissipation pathways theory to the Fenna-Matthews-Olson (FMO) complex. Additional information in the Supporting Information includes the electronic Hamiltonian for the FMO complex, the spectral densities used, the methodology for extracting the vibrational modes, and the dissipation dynamics for a different initial condition. We also present a set of animations for the extracted vibrational modes around 200 cm-1 in each bacteriochlorophyll.